OpenStudy (anonymous):

find dy/dx of y=pi^[(sin e^x)/(1-cos 2x)]?

OpenStudy (anonymous):

(ln(pi))pi^sine^x/1-cos2x times e^xcose^x(1-cos2x)-(sine^x)(2cos2x)/(1-cos2x)^2

OpenStudy (vijay):

when y= 2^x dy/dx = 2^(x) ln(2) think like this way, u may find the answer, also think pi is a number (constant)

OpenStudy (anonymous):

yeah that the rule but you also need to multiply by dy/dx of x